# Soft gravitons as Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter spacetimes

### v2

@@ Abstract
I demonstrate that soft graviton modes in de Sitter spacetimes are the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. I then show that any local measurement, including the effects of the environment, will collapse the symmetric state onto the broken state in the large volume limit. In any discussion involving observers, de Sitter spacetimes are, therefore, best described globally by the broken phase, while local observers, in the small volume limit, can not discriminate between different degenerate global vacuum states and are therefore best described by the symmetric state. As a consequence, a small Hubbled-sized local region initially in the symmetric state will, after a time-scale corresponding to the Page time of de Sitter space, have expanded to a large region in the broken state. This illuminates the physical nature of soft graviton modes in de Sitter spacetimes.
@@ I Introduction
It was argued some years ago in the context of black holes that the Goldstone modes of the spontaneously broken asymptotic symmetries of the black holes, the Bondi-Metzner-Sachs (BMS) supertranslations, are the soft graviton modes satisfying Weinberg’s soft graviton theorems He:2014laa ; Hawking:2016msc . There is a very analogous story in de Sitter space. The equivalent of the Weinberg soft theorem satisfied by the Weinberg adiabatic soft graviton modes is the Maldacena consistency relations, and the asymptotic symmetries of de Sitter space are the diffeomorphisms in three-dimensional Euclidean space Anninos:2010zf . Here I will show that the soft graviton modes are, indeed, also the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. In cosmology, we also think of soft graviton modes in de Sitter spacetimes as frozen-out super-horizon gravitational fluctuations, described by the traceless and transverse spatial component of the graviton.
In order to show that the superhorizon soft graviton is the Goldstone of the spontaneously broken asymptotic symmetry group of de Sitter space, I will discuss the symmetry-breaking pattern and then show that the counting of Goldstone modes adds up correctly. At the linearized level, the anisotropic scaling symmetry generated by symmetric traceless matrices is broken, leaving five broken generators. However, in the case of broken space-time symmetries, it is well known that Goldstone’s theorem for internal symmetries does not trivially apply. As in the case with spontaneously broken non-Lorentz invariant symmetries Nielsen:1975hm , also in the case of space-time symmetries, the number of massless Goldstone modes can be different than the number of broken symmetry generators. In the case of unbroken translations, local translation in the unbroken directions can be used to relate the Goldstone modes to each other Ivanov:1975zq ; Low:2001bw . With three unbroken translations, only two physical Goldstone modes remain out of the five broken generators. As I will show below, these are the two massless Goldstone modes of the soft graviton, which remain as physical modes. This argument generalizes to any dimension.
@@ II Spontaneously broken asymptotic symmetries and soft modes as Goldstones
In $n$ Euclidian dimensions the general covariance implies that the index, $j$, up or down transforms as covariant or contravariant vectors under the linear matrix multiplication of matrices given by $\partial x^{i}/\partial\tilde{x}^{j}$ $\partial\tilde{x}^{j}/\partial x^{i}$ at the level of linearized infinitesimal coordinate transformation. The linearized coordinate transformations are elements of the symmetry group $GL(n,\mathbb{R})$ with dimension $n^{2}$. $GL(n,\mathbb{R})$ can be written as $SL(n,\mathbb{R})\times\mathbb{R}$, where $SL(n,\mathbb{R})$ has $SO(n)$ as a subgroup with dimension $n(n-1)/2$ generated by the antisymmetric tensors $J_{ij}$, while the other $n(n+1)/2-1$ symmetric traceless generators $S_{ij}$ generate shear deformations. The dilatation, $D$, is part of $GL(n,\mathbb{R})$, but not $SL(n,\mathbb{R})$, which has dimension $n^{2}-1$. Including translation invariance enhances $GL(n,\mathbb{R})$ to the affine group $Aff(n,R))=GL(n,\mathbb{R})\times P_{n}$, where $P_{n}$ is the group of translations. The affine group, therefore, has the Poincaré group as a subgroup. The full diffeomorphism group is the closure of the affine group and the special conformal group. Any generator of the full diffeomorphism group can be written in terms of (infinite) combinations of the generators of $SL(n,\mathbb{R})$, as well as the generators of translations $P_{n}$, dilatations $D$ and special conformal transformations $K_{n}$ Borisov:1974bn . Important for the discussion below, at the linearized level, the generators of the asymptotic symmetry group are therefore $S_{ij},J_{ij},P_{i},K_{i},D$.
When we restrict ourselves to pure de Sitter space (with unbroken time translation), the relevant gravitational degrees of freedom are contained in the transverse and traceless graviton $\gamma_{ij}$, which be extracted from the metric in the planar slicing
@@
We have ignored the lapse and shift, which decay, and therefore are unimportant on the asymptotic boundary. The asymptotic symmetry transformations act on this metric as 1203.6351
@@
Let us consider a frozen-in soft graviton mode. It is the zero momentum limit of a plane wave model with momentum $k^{L}\to 0$. We can think of this mode as being a symmetric $3\times 3$ dimensional traceless matrix $\gamma_{ij}^{L}$ This mode is generated by a linear diffeomorphism of the form
@@
This is an equation that specifies the three functions $c_{i}(x)$, and therefore, we have only $5-3=2$ real Goldstone modes. This is an example of the fact that when spacetime symmetries are involved, the usual rule that the number of Goldstones of the breaking $G\supset H$ is the dimension of the coset $G/H$ does not work.
In fact, this argument works in any dimension. In $D=1+n$ spacetime dimensions, the $SL(n,\mathbb{R})$ of the asymptotic symmetry group is broken down to $SO(n,\mathbb{R})$, leaving $n(n+1)/2-1$ broken shear transformations. Now out of those $n$ can be related by unbroken spatial translations, leaving $n(n-1)/2-1=D(D-3)/2$ Goldstone modes, which is the number of degrees of freedom of a graviton in $D=1+n$ spacetime dimensions.
@@
In general, for a spontaneously broken symmetry generated by the charge, $Q$, there exists an order parameter $[Q,\mathcal{O}]=-i\delta\mathcal{O}$ for some operator $\mathcal{O}$ Brauner:2010wm . Using now $\xi=\xi_{i}^{L}$ given by Eq.(9) in Eq.(13), and substituting the expression into the standard expression for the Nöther charge
@@
We saw above that rotational invariance is unbroken at zeroth order in $\gamma$. However, since the order parameter is not invariant under rotations, it breaks rotational symmetry. Rotational symmetry is, therefore, broken at linear order in $\gamma$. In fact, this can be seen from taking $\xi_{i}=-\omega_{ij}x^{j}$, with $\omega_{ij}=-\omega_{ji}$ corresponding to a rotation. In this case, we have, when the rotation acts on $\gamma_{ij}$ at higher order, $\delta\gamma_{ij}=\omega_{jl}\gamma_{li}+\omega_{il}\gamma_{lj}$, and so $\partial_{i}\delta\gamma_{ij}=\partial_{i}\omega_{il}\gamma_{lj}\neq 0$, such that $\delta\gamma_{ij}$ generated by a rotation acting on $\gamma_{ij}(x)$ is not transverse, meaning that the rotations leave the transverse and traceless gauge at linear order in $\gamma$. Thus, at second order, rotational symmetry is nonlinearly realized on the Goldstone mode. By using that the angular momentum operator takes the form $J^{i}\propto\int d^{3}x\epsilon^{ijk}\gamma_{mk}\dot{\gamma}_{ml}$, one also finds $[J^{i},\gamma_{ij}]\neq 0$, as expected for the generator of a broken symmetry. So, while the anisotropic scaling symmetries, the shear transformations, are spontaneously broken at zeroth order in $\gamma$, their breaking induces a breaking of rotational symmetry at higher order.
@@
The Goldstone modes are the long-range spatial fluctuations of the order parameter $\gamma_{ij}^{L}$ Thus, the soft modes are Goldstones of the spontaneously broken asymptotic symmetries. This also means that the Lagrangian for the Goldstone modes is of the general form Naegels:2021ivf
@@ III Symmetry breaking: local versus global observers
[…]
@@
$0^{\prime}\right|\gamma_{ij}\left|0^{\prime}\right>$$\equiv$ $\gamma_{L}\right|\gamma_{ij}\left|\gamma_{L}\right>=\left<0\right|e^{-iQ_{{\gamma}_{L}}}\gamma_{ij}e^{iQ_{{\gamma}_{L}}}\left|0\right>$ (20)
@@
assuming that $0\right|\gamma_{ij}\left|0\right>=0$, and defining $Q(\xi_{i}^{L})\equiv Q_{{\gamma}_{L}}$. Thus, we see that the vacuum is infinitely degenerate, with each of the $\left|\gamma_{L}\right>$ states, related by symmetry transformations, representing different vacuum states of spontaneously broken anisotropic scaling symmetry with different values of $\gamma_{ij}^{L}$.
@@
The point is that if one takes the ensemble average of $\gamma_{ij}\right>$ in many local vacuum states, one will find that it vanishes. Therefore the local state is well represented by the symmetric state, $\left|\Psi_{\text{sym}}\right>$ which is a superposition of all possible values of $\gamma_{L}$. In this local state, $\gamma_{L}$ has no physical meaning, as the state does not depend on it. On the other hand, when comparing $\gamma_{ij}\right>$ in different local patches, one will see that they correlate in the infinite volume limit. The infinite volume is, therefore, well described by a single choice of $\gamma_{ij}^{L}(x)$ and the symmetry is broken by the global state, which is well represented by $\left|\Psi_{\gamma_{L}}\right>$. It is the act of measurement, or equivalently, the environment, which breaks the symmetry. Thus, once we include observers in the discussion of de Sitter space, the symmetry is spontaneously broken. This emphasizes the old mantra that in any discussion of a quantum system, the observers are part of the system bohr .
@@ VI Conclusions and outlook
Through examples, I discussed how this is just another formulation of the usual consistency relations and can be viewed as the de Sitter version of the equivalence between soft-theorems and asymptotic symmetries, which, when also thought of as a gravitational memory effect Ferreira:2017ogo , completes the triangle relating asymptotic symmetries, soft theorems and gravitational memory Strominger:2017zoo .
There are different applications and extensions of my work one could consider. Extending the Lagrangian of the Goldstone mode to higher orders, one has a different approach to an effective field theory of inflation or dark energy. In this context, it could also be interesting to extend our formalism to modified gravity theories or massive gravity theories. In addition, it is know that the charge, $Q$, has a dual Conformal Field Theory (CFT) interpretation as a topological charge Kehagias:2017rpe , and it would be interesting to explore the CFT dual of the spontaneous symmetry breaking. Finally, it would be interesting if a de Sitter liquid crystal, with the same symmetry-breaking pattern as discussed here, could be realized in the laboratory, as an analog system of de Sitter spacetimes and eternal inflation.

### v3

@@ Abstract
I demonstrate that soft graviton modes in de Sitter spacetimes are the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. I then show that any local measurement, including the effects of the environment, will collapse the symmetric state onto the broken state in the large volume limit. In any discussion involving observers, de Sitter spacetimes are, therefore, best described globally by the broken phase, while local observers, in the small volume limit, can not discriminate between different degenerate global vacuum states and are therefore best described by the symmetric state. As a consequence, a small Hubble-sized local region initially in the symmetric state will, after a time scale corresponding to the Page time of de Sitter space, have expanded to a large region in the broken state. This illuminates the physical nature of soft graviton modes in de Sitter spacetimes.
@@ I Introduction
It was argued some years ago in the context of black holes that the Goldstone modes of the spontaneously broken asymptotic symmetries of the black holes, the Bondi-Metzner-Sachs (BMS) supertranslations, are the soft graviton modes satisfying Weinberg’s soft graviton theorems He:2014laa ; Hawking:2016msc . There is a very analogous story in de Sitter space. The equivalent of the Weinberg soft theorem satisfied by the Weinberg adiabatic soft graviton modes is the Maldacena consistency relations, and the asymptotic symmetries of de Sitter space are the diffeomorphisms in three-dimensional Euclidean space Anninos:2010zf . Here, I will show that the soft graviton modes are, indeed, also the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. In cosmology, we also think of soft graviton modes in de Sitter spacetimes as frozen-out super-horizon gravitational fluctuations, described by the traceless and transverse spatial component of the graviton.
In order to show that the superhorizon soft graviton is the Goldstone of the spontaneously broken asymptotic symmetry group of de Sitter space, I will discuss the symmetry-breaking pattern and then show that the counting of Goldstone modes adds up correctly. At the linearized level, the anisotropic scaling symmetry generated by symmetric traceless matrices is broken, leaving five broken generators. However, in the case of broken space-time symmetries, it is well known that Goldstone’s theorem for internal symmetries does not trivially apply. As in the case with spontaneously broken non-Lorentz invariant symmetries Nielsen:1975hm , also in the case of space-time symmetries, the number of massless Goldstone modes can be different from the number of broken symmetry generators. In the case of unbroken translations, local translation in the unbroken directions can be used to relate the Goldstone modes to each other Ivanov:1975zq ; Low:2001bw . With three unbroken translations, only two physical Goldstone modes remain out of the five broken generators. As I will show below, these are the two massless Goldstone modes of the soft graviton, which remain as physical modes. This argument generalizes to any dimension.
@@ II Spontaneously broken asymptotic symmetries and soft modes as Goldstones
In $n$ Euclidian dimensions the general covariance implies that the index, $j$, up or down transforms as covariant or contravariant vectors under the linear matrix multiplication of matrices given by $\partial x^{i}/\partial\tilde{x}^{j}$ $\partial\tilde{x}^{j}/\partial x^{i}$ at the level of linearized infinitesimal coordinate transformation. The linearized coordinate transformations are elements of the symmetry group $GL(n,\mathbb{R})$ with dimension $n^{2}$. $GL(n,\mathbb{R})$ can be written as $SL(n,\mathbb{R})\times\mathbb{R}$, where $SL(n,\mathbb{R})$ has $SO(n)$ as a subgroup with dimension n(n1)/2n(n-1)/2 generated by the antisymmetric tensors $J_{ij}$, while the other n(n+1)/21n(n+1)/2-1 symmetric traceless generators $S_{ij}$ generate shear deformations. The dilatation, $D$, is part of $GL(n,\mathbb{R})$, but not $SL(n,\mathbb{R})$, which has dimension $n^{2}-1$. Including translation invariance enhances $GL(n,\mathbb{R})$ to the affine group $Aff(n,R))=GL(n,\mathbb{R})\times P_{n}$, where $P_{n}$ is the group of translations. The affine group, therefore, has the Poincaré group as a subgroup. The full diffeomorphism group is the closure of the affine group and the special conformal group. Any generator of the full diffeomorphism group can be written in terms of (infinite) combinations of the generators of $SL(n,\mathbb{R})$, as well as the generators of translations $P_{n}$, dilatations $D$ and special conformal transformations $K_{n}$ Borisov:1974bn . Important for the discussion below, at the linearized level, the generators of the asymptotic symmetry group are therefore $S_{ij},J_{ij},P_{i},K_{i},D$.
When we restrict ourselves to pure de Sitter space (with unbroken time translation), the relevant gravitational degrees of freedom are contained in the transverse and traceless graviton $\gamma_{ij}$, which can be extracted from the metric in the planar slicing
@@
We have ignored the lapse and shift, which decay and, therefore, are unimportant on the asymptotic boundary. The asymptotic symmetry transformations act on this metric as 1203.6351
@@
Let us consider a frozen-in soft graviton mode. It is the zero momentum limit of a plane wave model with momentum $k^{L}\to 0$. We can think of this mode as being a symmetric $3\times 3$ dimensional traceless matrix $\gamma_{ij}^{L}$ . This mode is generated by a linear diffeomorphism of the form
@@
This is an equation that specifies the three functions $c_{i}(x)$, and therefore, we have only 53=25-3=2 real Goldstone modes. This is an example of the fact that when spacetime symmetries are involved, the usual rule that the number of Goldstones of the breaking $G\supset H$ is the dimension of the coset $G/H$ does not work.
In fact, this argument works in any dimension. In $D=1+n$ spacetime dimensions, the $SL(n,\mathbb{R})$ of the asymptotic symmetry group is broken down to $SO(n,\mathbb{R})$, leaving n(n+1)/21n(n+1)/2-1 broken shear transformations. Now out of those $n$ can be related by unbroken spatial translations, leaving n(n1)/21=D(D3)/2n(n-1)/2-1=D(D-3)/2 Goldstone modes, which is the number of degrees of freedom of a graviton in $D=1+n$ spacetime dimensions.
@@
In general, for a spontaneously broken symmetry generated by the charge, $Q$, there exists an order parameter $Q,\mathcal{O}]=-i\delta\mathcal{O}$ for some operator $\mathcal{O}$ Brauner:2010wm . Using now $\xi=\xi_{i}^{L}$ given by Eq.(9) in Eq.(13), and substituting the expression into the standard expression for the Nöther charge
@@
We saw above that rotational invariance is unbroken at zeroth order in $\gamma$. However, since the order parameter is not invariant under rotations, it breaks rotational symmetry. Rotational symmetry is, therefore, broken at linear order in $\gamma$. In fact, this can be seen from taking $\xi_{i}=-\omega_{ij}x^{j}$, with $\omega_{ij}=-\omega_{ji}$ corresponding to a rotation. In this case, we have, when the rotation acts on $\gamma_{ij}$ at higher order, $\delta\gamma_{ij}=\omega_{jl}\gamma_{li}+\omega_{il}\gamma_{lj}$, and so $\partial_{i}\delta\gamma_{ij}=\partial_{i}\omega_{il}\gamma_{lj}\neq 0$, such that $\delta\gamma_{ij}$ generated by a rotation acting on $\gamma_{ij}(x)$ is not transverse, meaning that the rotations leave the transverse and traceless gauge at linear order in $\gamma$. Thus, at second order, rotational symmetry is nonlinearly realized on the Goldstone mode. By using the fact that the angular momentum operator takes the form $J^{i}\propto\int d^{3}x\epsilon^{ijk}\gamma_{mk}\dot{\gamma}_{ml}$, one also finds $J^{i},\gamma_{ij}]\neq 0$, as expected for the generator of a broken symmetry. So, while the anisotropic scaling symmetries, the shear transformations, are spontaneously broken at zeroth order in $\gamma$, their breaking induces a breaking of rotational symmetry at higher order.
@@
The Goldstone modes are the long-range spatial fluctuations of the order parameter $\gamma_{ij}^{L}$ . Thus, the soft modes are Goldstones of the spontaneously broken asymptotic symmetries. This also means that the Lagrangian for the Goldstone modes is of the general form Naegels:2021ivf
@@ III Symmetry breaking: local versus global observers
$\left|0^{\prime}\right>=e^{iQ_{\gamma_{L}}}\left|0\right>\equiv\left|\gamma_{L}\right$> (19)
@@
$\left<0^{\prime}\right|\gamma_{ij}\left|0^{\prime}\right$> $\equiv$ $\left<\gamma_{L}\right|\gamma_{ij}\left|\gamma_{L}\right>=\left<0\right|e^{-iQ_{{\gamma}_{L}}}\gamma_{ij}e^{iQ_{{\gamma}_{L}}}\left|0\right$> (20)
@@
assuming that $\left<0\right|\gamma_{ij}\left|0\right>=0$, and defining $Q(\xi_{i}^{L})\equiv Q_{{\gamma}_{L}}$. Thus, we see that the vacuum is infinitely degenerate, with each of the $\left|\gamma_{L}\right$> states, related by symmetry transformations, representing different vacuum states of spontaneously broken anisotropic scaling symmetry with different values of $\gamma_{ij}^{L}$.
@@
The point is that if one takes the ensemble average of $\left<\gamma_{ij}\right$> in many local vacuum states, one will find that it vanishes. Therefore, the local state is well represented by the symmetric state, $\left|\Psi_{\text{sym}}\right$>, which is a superposition of all possible values of $\gamma_{L}$. In this local state, $\gamma_{L}$ has no physical meaning, as the state does not depend on it. On the other hand, when comparing $\left<\gamma_{ij}\right$> in different local patches, one will see that they correlate in the infinite volume limit. The infinite volume is, therefore, well described by a single choice of $\gamma_{ij}^{L}(x)$ and the symmetry is broken by the global state, which is well represented by $\left|\Psi_{\gamma_{L}}\right$>. It is the act of measurement, or equivalently, the environment, which breaks the symmetry. Thus, once we include observers in the discussion of de Sitter space, the symmetry is spontaneously broken. This emphasizes the old mantra that in any discussion of a quantum system, the observers are part of the system bohr .
@@ VI Conclusions and outlook
Through examples, I discussed how this is just another formulation of the usual consistency relations and can be viewed as the de Sitter version of the equivalence between soft theorems and asymptotic symmetries, which, when also thought of as a gravitational memory effect Ferreira:2017ogo , completes the triangle relating asymptotic symmetries, soft theorems and gravitational memory Strominger:2017zoo .
There are different applications and extensions of my work that one could consider. Extending the Lagrangian of the Goldstone mode to higher orders, one has a different approach to an effective field theory of inflation or dark energy. In this context, it could also be interesting to extend our formalism to modified gravity theories or massive gravity theories. In addition, it is known that the charge, $Q$, has a dual Conformal Field Theory (CFT) interpretation as a topological charge Kehagias:2017rpe , and it would be interesting to explore the CFT dual of the spontaneous symmetry breaking. Finally, it would be interesting if a de Sitter liquid crystal, with the same symmetry-breaking pattern as discussed here, could be realized in the laboratory, as an analogue system of de Sitter spacetimes and eternal inflation.